OpenStudy (anonymous):

What is the antiderivative of 2t^2*sqrt t+1/(t*sqrt t)

OpenStudy (anonymous):

Do you mean the antiderivative of the following?\[2t ^{2}\sqrt{t}+\frac{ 1 }{ t \sqrt{t} }\]What difficulty are you having? How far have you got with it? Remember the following:-\[\sqrt{x}=x ^{\frac{ 1 }{ 2 }} \text{(Nonnegative values only.)}\]\[x^{a} x^{b}=x^{a+b}\]\[d(u+v)=du+dv\]

OpenStudy (anonymous):

I suggest you to use sqrt(2) instead of sqrt 2 for us to better understand your formulae. Could you restate your question using that notation? Cheers!

OpenStudy (anonymous):

IT WOULD BE-(4(T^3.5)/7 )-2T^(-0.5) +C

OpenStudy (anonymous):